OpenStudy (anonymous):

Choose all answers that are correct. 1. (1 pt) For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction @pgpilot326 @Destinymasha @shrutipande9 @e.mccormick

OpenStudy (anonymous):

I dont really know. :/

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

0

OpenStudy (anonymous):

take and perform the operations (one at a time) with all possible combinations of elements and see if the results are conatined within the original set. if so, then the set is closed under that operation. if not, then the set is not closed under that operation. for example, let's check addition. we have 2 elements. 0+0= 0 (in the original set) 0+1= 1 (in the original set) 1+0= 1 (in the original set) 1+1= 2 (not in the original set) => not closed under addition (because 2 is not in the original set) does that make sense? you have to do the rest but I'll check your work (if you like)

OpenStudy (anonymous):

hello, you there?